Hello, I am hoping with some help on thisequation and what I have so far.
Here it is: 2(x-3)-3(x+5)=3(x-2)-7
what I have come up with so far is this: (2x-2*3)-(3x+3*5)=(3*-3*2)-7 with the next step being this: (2x-6)-(3x+15)=(3x-6)-7
Anyways I end up with the final answer being X=-2 for this equation am I right on this one? Your input would be greatly appreciated.
Thank You, Barb Neely

Answer by mathnard(10)

yes, x=-2. You can always check your answer by plugging it back in.

2((-2)-3) - 3((-2)+5) = 3((-2)-2) - 7
which simplifies to
2(-5) - 3(3) = 3(-4) - 7
-10-9=-12-7
-19=-19
Which is true!!

So we have checked the solution and know it is correct!